Write the equation of the line that passes through (-4,7) with a slope of a half.

thanks

Answer by nerdybill(7384)

Write the equation of the line that passes through (-4,7) with a slope of a half.
.
Given:
point (-4,7)
slope, m, 1/2
.
Plug it all into the "point slope" form:
y-y1 = m(x-x1)
y-7 = (1/2)(x-(-4))
y-7 = (1/2)(x+4)
y-7 = (1/2)x + 2
y = (1/2)x + 9 (this is the slope-intercept form:what they're looking for)